OpenStudy (anonymous):

Solve for x. x/5 = 20

OpenStudy (♪chibiterasu):

x/5 = 20 x/5 • 5 = 20 • 5 x = 100

OpenStudy (austinl):

\(\dfrac{x}{5}=20\) Just multiply both sides by 5. \(5(\dfrac{x}{5})=5(20)\) \(x=~?\)

OpenStudy (anonymous):

x/5=20 move 5 to the other side multiplying it with 20 x=20*5 x=100

OpenStudy (♪chibiterasu):

Too late :P